A gas has a volume of 255 mL, a pressure of 733 mm Hg, a temperature of 548 ℃, and a mass of 1.22g. What is the molar mass? using pV = nRT
1
Expert's answer
2016-01-25T07:22:56-0500
V = 255 mL = 0,255*10–3 m3
p = 733 mm Hg = 97 725 Pa
T = 548 ℃ = 821 К
m = 1,22 g = 1,22*10–3 kg
R = 8,314

n = m/M; M = mRT/pV
M = 1,22*10–3 *8,314*821/97 725*0,255*10–3 = 0,334 kg/mol